Problem 1.6: A piston-cylinder device contains 0.05 m3 of a gas initially


at 200 MPa. At this state, a linear spring that has spring constant of 150

kN/m is touching the piston but exerting no force on it. Now heat is

transferred to the gas, causing the piston to rise and compress the spring until the volume inside the cylinder doubles. If the cross sectional area of the piston is 0.25 m2, determine the fraction of workdone by the gas against

the spring to compress it.





1
Expert's answer
2021-07-20T15:23:09-0400

Gives

K=150KN/m


"x_2=0.2,x_1=0,V=0.05m^3P_1=200kpa,A=0.25m^2"

We know that

"W(s)=\frac{1}{2}k(x_2^2-x_1^2)"

Put value


"W(s)=\frac{1}{2}\times150\times(0.2^2-0)KJ"

"W(s)=3KJ"